# How to Understand Whole Software Repository?

### v1

@@ 3.3. MCTS-Enhanced Repository Understanding
Inspired by the success of large language models in in-context learning (ICL) (Dong et al., 2022; Work, [n. d.]) and Chain-of-Thought (CoT) (Wei et al., 2022) methods, we propose a reward method based on ICL and CoT, aiming to provide a more accurate and reliable solution for relevance evaluation in software repository. Our approach leverages the advanced ability of large language models to learn and optimize reward functions from limited examples of programming practice to accurately assess the correlation between leaf nodes and problem descriptions. Specifically, we first use ICL to let the language model learn to understand the core functions and operating modes of the reward function in a given context. Then, the CoT is used to enable the model to conduct in-depth reasoning based on the specific information in the question and code snippets to evaluate the correlation of leaf nodes. The reward function prompt template we designed (see Figure 2) starts with a guided system prompt that clearly points out the goals and responsibilities of the reward function. Then, through a series of example combinations of ¡issue description, code snippets, thinking process, results¿, the input, output and reasoning chain in the scoring process are demonstrated. Finally, the prompt ends with a new set of issue descriptions and code snippets, at which point the model is expected to learn the intermediate reasoning steps from the given examples and output corresponding reward scores. Finally, we only keep the nodes with a reward score of no less than 6 and return their content and structural dependencies.
@@ 3.4. Information Utilization & Patch Generation
At this stage, RepoUnderstander first summarizes the whole repository experience, then obtains code snippet information dynamicly on this basis, and finally generates patches that try to solve the problem. The three steps are detailed below.
Repository Summary. To more effectively utilize the global repository information collected during the repository understanding phase, we introduce a summary agent. The agent aims to systematically analyze and summarize the code snippets collected in the repository knowledge graph and submitted issues , and then plan how to solve the problem, thereby forming an experience of the entire repository. Specifically, the summary agent takes the issue and the collected relevant code fragments as input, and then outputs a summary of the relevant fragments in sequence and plans a solution. The specific prompt template is shown in Figure 3. Since the collected global repository information may be complex and contain a large number of code fragments and annotation descriptions, we only use the location description of the relevant code fragments (i.e., structural dependencies in the repository) and the output of the Summary Agent (i.e., summary and planning) as RepoUnderstander’s experience to guide subsequent actions. This experience does not include specific function implementation, but only focuses on overall repository experience guidance. The location description is formalized as ¡file¿a.py¡/file¿¡class¿Class A¡/class¿¡func¿func a¡/func¿, and the summary agent output is as shown in Figure 3.
@@
Dynamic information acquisition. Global experience information is RepoUnderstander’s experience summary of relevant information in the current workspace, which can help the language model understand issues and find solutions more quickly. In the process of solving problems, in order to make full use of this global experience information, RepoUnderstander futher needs to dynamically extract local information from the current repository, including specific classes, functions and code snippets in the repository.
The ReAct (Yao et al., 2022) framework (i.e., Reson then Act) guides the model to generate inference trajectories and task-specific actions in a staggered manner, allowing the model to interact with the code repository and collect information. Specifically, the ReAct framework first generates reasoning paths through the Chain-of-Thought (Wei et al., 2022), and then outputs actual actions based on the reasoning results. Therefore, by using ReAct method, RepoUnderstander can call the corresponding search API according to task requirements and dynamically extract local information from the current repository to collect relevant context. We follow AutoCodeRover’s search API method (Zhang et al., 2024b), using the three-layer search method of search_class, search_method, and search_code. Specifically, RepoUnderstander first independently determines the API that needs to be called. Then the retrieval API will search for classes, methods and code snippets in the repository knowledge graph, and finally return the results to the agent.
Patch Generation. In the patch generation step, RepoUnderstander first locates faults based on the summary of global experience and dynamic information, extracts the context of code snippets that may need to be modified, and then generates modified code snippets. Finally, a diff is generated based on the code snippet before modification and the code snippet after modification, and is returned as the final result. If a diff is incorrect due to syntax, we will retry until an applicable patch with correct syntax is generated. We follow AutoCodeRover (Zhang et al., 2024b) and set the maximum number of retries to 3 to ensure that the generated patch can be applied as much as possible.
@@ 4. Experiment
To validate the performance of RepoUnderstander, we conduct experiments and compare it with other LLMs and agents to demonstrate its superiority (§4.2, §4.4). In addition, we found that there is an information asymmetry problem in SWE-bench (Jimenez et al., 2024) caused by new added function names and variable names in the test patch (§4.3). We proposed issue patches to new feature types issues to solve this problem and test agents on the new FIX version test set. Finally, we systematically analyzed the problem-solving capabilities of RepoUnderstander at different stages (§4.5).
@@ 4.1. Experimental Setup
Baselines. We compare RepoUnderstander with two types of baselines. The first category is the RAG baselines (Jimenez et al., 2024). This type of baseline uses the BM25 method to retrieve code base files related to the issue and inputs them into LLM to directly generate patch files that solve the problem. The second type of baseline is the agents baseline (i.e., AutoCodeRover (Zhang et al., 2024b) and SWE-agent (Yang et al., [n. d.])), which locates the problem through complex multiple rounds of interaction and execution feedback, and finally generates a patch to solve the problem through iterative verification.
Metrics. Following the SWE-bench (Jimenez et al., 2024), We evaluate the effectiveness of RepoUnderstander, using the percentage of resolved instances and the patch application rate. Among them, the patch application rate refers to the proportion of instances where code changes are successfully generated and can be applied to existing code bases using Git tools. Resolved ratio represents the overall effectiveness of solving actual GitHub issues, and application ratio reflects the intermediate results of patch availability.
Configurations. All results, ablations, and result analyzes of RepoUnderstander use the GPT4-Turbo model (i.e., gpt-4-1106-preview (Achiam et al., 2023), the same model with SWE-agent (Yang et al., [n. d.])). We use ast [^2] https://docs.python.org/3/library/ast.html and Jedi [^3] https://github.com/davidhalter/jedi library to parse repository and obtain syntax structures and dependencies of repository. In MCTS-Enhanced Repository Understanding stage, we set the number of search iterations to 600 and maximum search time to 300 seconds. In information Utilization & Patch Generation stage, we set the maximun number of summary code snippets to 10. SWE-bench has a relatively complex environment configuration. Thanks to the development of the open source community, we use the well-build open source docker of the AutoCodeRover (Zhang et al., 2024b) team for experiments.
@@ 4.2. Comparison Experiment
Method Resolved Apply
@@
SWE-Llama 7B 1.33% (4) 38.00%
SWE-Llama 13B 1.00% (3) 38.00%
ChatGPT-3.5 0.33% (1) 10.33%
GPT-4 2.67% (8) 29.67%
Claude-2 3.00% (9) 33.00%
Claude-3 Opus 4.33% (13) 51.67%
@@
AutoCodeRover 16.11% (48) 83.00%
SWE-agent 18.00% (54) 93.00%
RepoUnderstander 21.33% (64) 85.67%
@@
Table 1. Main results for RepoUnderstander performance on the SWE-bench-lite test set. The numbers in brackets indicate the number of issues solved.
We first evaluate the effectiveness of RepoUnderstander in SWE-bench Lite (300 instances). The performance comparison analysis between RepoUnderstander and other methods is shown in Table 1. In each instance, we provide a natural language description from a real-world software engineering problem and a local code repository of corresponding versions, asking the model to solve the problem and generate patches that can pass local automated testing. Resolved reflects the end-to-end ability of the current RAG LLM system and Agent system to solve software engineering problems. The results show that RepoUnderstander is significantly better than other RAG and Agent systems, achieving SOTA performance on the test set. Compared with the RAG system, our method improves performance by nearly 5 times. Compared with the state-of-the-art Agent system, we improve the accuracy of SWE-agent by 18.5%. These excellent performances demonstrate the advancement of our approach. In addition, the Apply application rate indicates the availability of generated patches. We found that Agent-based systems all achieved high availability, while RAG-based systems have lower availability, which proves that agent systems may be an important means to automatically solve software engineering tasks. SWE-agent has the highest Apply application rate due to the introduction of its running feedback capability, which shows that running feedback is an effective way. This paper focuses more on the understanding of the entire repository information. We will integrate running feedback in future work.
In addition, we also compared the issue-solving distribution diagrams of three Agent-based methods, and the results are shown in Figure 4. We found that our method is very complementary to the SWE-agent method. The two methods jointly solved 80 examples, achieving a task resolved rate of 26.67% (see Table 2), which further illustrates the complementarity of our method and the execution feedback method. We will provide a detail discussion and combination of the two methods in future work.
@@
Figure 4. Venn diagrams of resolved cases of RepoUnderstander, SWE-agent and AutoCodeRover.
Method Resolved Apply
ACR & SWE-agent 24.33% (73) 98.00%
RepoU & ACR 25.33% (76) 94.67%
RepoU & SWE-agent 26.67% (80) 99.67%
Table 2. Venn diagram analysis of our method and baselines.
4.3. Dataset Analysis & Fix
User issues usually include bug reports, feature requests, and enhancements, etc (Moran Altarac, 2024; Merten et al., 2016; Krishna et al., 2018; wiki koha, 2024). In SWE-bench dataset (Jimenez et al., 2024), we found that there is information asymmetry for issues of the feature request type, such as adding functions or adding parameter definitions. Specifically, since the test patch contains the signature information of the new features, and the LLM Agents input lacks this interface specification information, the agent model may not be able to correctly understand the full context of the problem. Even if the logic of the generated patch is correct, errors may occur during testing. This information asymmetry may affect the performance evaluation and practical application of the LLM agents. From the perspective of software engineering practice, new features are usually defined and specified in the system design document. In actual development, the information of these new features should be specified rather than inferred by the agent model. Therefore, merging the new feature interface specification in the test patch into the problem description can enable the agent model to better understand the problem and reduce the impact of information asymmetry.
We made a fix for the SWE-bench Lite dataset and proposed a FIX version. Specifically, we integrated the new feature interface specifications in the test patch into the problem description through manual analysis, and guided the agent model to generate patches based on the complete problem description. In total, we fixed 45/300 instances in SWE-bench Lite. This method can better reflect the reality in practice, reduce the problems caused by information asymmetry, and thus improve the credibility and effectiveness of automatic software engineering technology in practical applications. As shown in Table 3, our experimental results show that on the FIX version test set, the methods RepoUnderstander and AutoCodeRover (Zhang et al., 2024b) have all improved, among which RepoUnderstander performs best, further demonstrating the superior performance of RepoUnderstander. We look forward to subsequent agent system work reporting their performance on the FIX version.
Method Resolved Apply
SWE-bench-Lite
AutoCodeRover 16.11% (48) 83.00%
RepoUnderstander 21.33% (64) 85.67%
SWE-bench-Lite-FIX
AutoCodeRover 18.00% (54) 84.00%
RepoUnderstander 23.00% (69) 88.00%
Table 3. Results for RepoUnderstander performance on the SWE-bench-Lite-FIX test set.
@@ 4.4.1. Module Analysis
This ablation experiment aims to study the effectiveness of RepoUnderstander’s global repository understanding component. (1) Remove MCTS & summary modules: RepoUnderstander has no prior knowledge of the repository structure and functions, that is, it lacks empirical information about the whole repository and can only locate relevant code snippets by searching through limited information in the issue. (2) Remove only the summary module: Only the signature and dependency structure of relevant information in the repository obtained by MCTS are used as global experience, and the summary and planning of information are removed. This experiment aims to verify the effectiveness of the summary agent, i.e., the importance of comprehensive summary of repository information. (3) Add a review agent module: After RepoUnderstander generates a patch that can be applied, in order to simulate the code review process in the development process, a static review of the patch by the review agent is added to discover possible defects in the newly generated code. If there is a defect, the patch is regenerated according to the review reason until a patch that passes the review is generated. This process is repeated up to three times.
Our experimental results demonstrate the importance of global experience and the effectiveness of the summary agent. As shown in Table 4, removing these modules all resulted in a drop in the performance of RepoUnderstander, especially after removing the MCTS & summary agent; the number of problem instances solved decreased from 64 to 48, which highlights the importance of global experience for automatically solving repository-level issues. In addition, we found that after adding the review agent, the performance of RepoUnderstander dropped, suggesting the limitations of static review. We speculate that the LLM-based static review may only rely on the surface grammatical information of the code and cannot fully understand the semantic meaning of the code. Therefore, the static review may ignore some hidden logical errors or illogical situations in the code. Therefore, we suggest that subsequent work can combine dynamic program analysis (Zhang et al., 2023a; Deng et al., 2023; Xia et al., 2024) such as program instrumentation (Hollingsworth et al., 1994; Huang, 1978) to improve the reliability of the LLM Agent.
@@
RepoUnderstander 21.33% (64) 85.67%
- w/o. summary 17.67% (53) 85.33%
- w/o. mcts & summary 16.00% (48) 80.33%
- w. review 18.33% (55) 87.67%
Table 4. Ablation results of RepoUnderstander.

### v2

@@ 3.3. MCTS-Enhanced Repository Understanding
Drawing upon previous work on in-context learning (ICL) and Chain-of-Thought (CoT) (Dong et al., 2022; Work, [n. d.]; Wei et al., 2022), we employ a reward method based on ICL and CoT to provide reward scores. Our approach leverages the advanced ability of LLMs to learn and optimize reward functions from limited examples of programming practice to accurately assess the correlation between leaf nodes and problem descriptions. Specifically, we first use ICL to let the language model learn to understand the core functions and operating modes of the reward function in a given context. Then, the CoT is used to enable the model to conduct in-depth reasoning based on the specific information in the question and code snippets to evaluate the correlation of leaf nodes. The reward function prompt template we designed (see Figure 2) starts with a guided system prompt that clearly points out the goals and responsibilities of the reward function. Then, through a series of example combinations of ¡issue description, code snippets, thinking process, results¿, the input, output and reasoning chain in the scoring process are demonstrated. Finally, the prompt ends with a new set of issue descriptions and code snippets, at which point the model is expected to learn the intermediate reasoning steps from the given examples and output corresponding reward scores. Finally, we only keep the nodes with a reward score of no less than 6 and return their content and structural dependencies.
@@ 3.4. Information Utilization & Patch Generation
At this stage, LingmaAgent first summarizes the whole repository experience, then obtains code snippet information dynamicly on this basis, and finally generates patches that try to solve the problem. The three steps are detailed below.
Repository Summary. To more effectively utilize the global repository information collected during the repository understanding phase, we introduce a summary agent. The agent aims to systematically analyze and summarize the code snippets collected in the repository knowledge graph and submitted issues , and then plan how to solve the problem, thereby forming an experience of the entire repository. Specifically, the summary agent takes the issue and the collected relevant code fragments as input, and then outputs a summary of the relevant fragments in sequence and plans a solution. The specific prompt template is shown in Figure 3. Since the collected global repository information may be complex and contain a large number of code fragments and annotation descriptions, we only use the location description of the relevant code fragments (i.e., structural dependencies in the repository) and the output of the Summary Agent (i.e., summary and planning) as LingmaAgent’s experience to guide subsequent actions. This experience does not include specific function implementation, but only focuses on overall repository experience guidance. The location description is formalized as ¡file¿a.py¡/file¿¡class¿Class A¡/class¿¡func¿func a¡/func¿, and the summary agent output is as shown in Figure 3.
@@
Dynamic information acquisition. Global experience information is LingmaAgent’s experience summary of relevant information in the current workspace, which can help the language model understand issues and find solutions more quickly. In the process of solving problems, in order to make full use of this global experience information, LingmaAgent futher needs to dynamically extract local information from the current repository, including specific classes, functions and code snippets in the repository.
The ReAct (Yao et al., 2022) framework (i.e., Reson then Act) guides the model to generate inference trajectories and task-specific actions in a staggered manner, allowing the model to interact with the code repository and collect information. Specifically, the ReAct framework first generates reasoning paths through the Chain-of-Thought (Wei et al., 2022), and then outputs actual actions based on the reasoning results. Therefore, by using ReAct method, LingmaAgent can call the corresponding search API according to task requirements and dynamically extract local information from the current repository to collect relevant context. We follow AutoCodeRover’s search API method (Zhang et al., 2024b), using the three-layer search method of search_class, search_method, and search_code. Specifically, LingmaAgent first independently determines the API that needs to be called. Then the retrieval API will search for classes, methods and code snippets in the repository knowledge graph, and finally return the results to the agent.
Patch Generation. In the patch generation step, LingmaAgent first locates faults based on the summary of global experience and dynamic information, extracts the context of code snippets that may need to be modified, and then generates modified code snippets. Finally, a diff is generated based on the code snippet before modification and the code snippet after modification, and is returned as the final result. If a diff is incorrect due to syntax, we will retry until an applicable patch with correct syntax is generated. We follow AutoCodeRover (Zhang et al., 2024b) and set the maximum number of retries to 3 to ensure that the generated patch can be applied as much as possible.
@@ 4. Experiment
To validate the performance of LingmaAgent, we conduct a series of comprehensive experiments and comparisons. We begin by comparing LingmaAgent with RAG-based and Agent-based systems on the SWE-bench Lite dataset (§4.2). We then assess the consistency and reliability of LingmaAgent’s performance across multiple runs (§4.3). In addition, we conduct detailed ablation studies to understand the contribution of each component in LingmaAgent (§4.4). Finally, we assess LingmaAgent’s effectiveness in industrial settings using an in-house dataset from Alibaba Cloud, testing both fully automated and human-in-the-loop scenarios (§4.5).
@@ 4.1. Experimental Setup
Baselines. We compare LingmaAgent with two types of baselines. The first category is the RAG baselines (Jimenez et al., 2024). This type of baseline uses the BM25 method to retrieve code base files related to the issue and inputs them into LLM to directly generate patch files that solve the problem. The second type of baseline is the agents baseline (i.e., AutoCodeRover (Zhang et al., 2024b) and SWE-agent (Yang et al., 2024)), which locates the problem through complex multiple rounds of interaction and execution feedback, and finally generates a patch to solve the problem through iterative verification.
Metrics. Following the SWE-bench (Jimenez et al., 2024), We evaluate the effectiveness of LingmaAgent, using the percentage of resolved instances and the patch application rate. Among them, the patch application rate refers to the proportion of instances where code changes are successfully generated and can be applied to existing code bases using Git tools. Resolved ratio represents the overall effectiveness of solving actual GitHub issues, and application ratio reflects the intermediate results of patch availability.
Configurations. All results, ablations, and result analyzes of LingmaAgent use the GPT4-Turbo model (i.e., gpt-4-1106-preview (Achiam et al., 2023), the same model with SWE-agent (Yang et al., 2024)). We use ast [^2] https://docs.python.org/3/library/ast.html and Jedi [^3] https://github.com/davidhalter/jedi library to parse repository and obtain syntax structures and dependencies of repository. In MCTS-Enhanced Repository Understanding stage, we set the number of search iterations to 600 and maximum search time to 300 seconds. In information Utilization & Patch Generation stage, we set the maximun number of summary code snippets to 10. SWE-bench has a relatively complex environment configuration. Thanks to the development of the open source community, we use the well-build open source docker of the AutoCodeRover (Zhang et al., 2024b) team for experiments.
@@ 4.2. Comparison Experiment
We first evaluate the effectiveness of LingmaAgent in SWE-bench Lite (300 instances). The performance comparison analysis between LingmaAgent and other methods is shown in Table 1. In each instance, we provide a natural language description from a real-world software engineering problem and a local code repository of corresponding versions, asking the model to solve the problem and generate patches that can pass local automated testing. Resolved reflects the end-to-end ability of the current RAG LLM system and Agent system to solve software engineering problems. The results show that LingmaAgent is significantly better than other RAG and Agent systems, achieving SOTA performance on the test set. Compared with the RAG system, our method improves performance by nearly 5 times. Compared with the state-of-the-art Agent system, we improve the accuracy of SWE-agent by 18.5%. These excellent performances demonstrate the advancement of our approach. In addition, the Apply application rate indicates the availability of generated patches. We found that Agent-based systems all achieved high availability, while RAG-based systems have lower availability, which proves that agent systems may be an important means to automatically solve software engineering tasks.
Method Resolved Apply Avg Cost
@@
SWE-Llama 7B 1.33% 38.00% -
SWE-Llama 13B 1.00% 38.00% -
GPT-4 2.67% 29.67% \$0.13
Claude-3 Opus 4.33% 51.67% \$0.25
@@
AutoCodeRover 16.11% 83.00% \$0.45
SWE-agent 18.00% 93.00% \$2.51
LingmaAgent 21.33% (18.5% $\uparrow$) 85.67% \$3.99
@@
Table 1. Main results for LingmaAgent performance on the SWE-bench-lite test set. The numbers in brackets indicate the number of issues solved.
Method Resolved Apply Avg Cost
ACR & SWE-agent 24.33% (73) 98.00% -
Lingma & ACR 25.33% (76) 94.67% -
Lingma & SWE-agent 26.67% (80) 99.67% -
LingmaAgent (w.feedback)
GPT-4 27.70% (83) 96.30% \$4.72
GPT-4o 28.33% (85) 95.33% \$2.39
Claude3.5 Sonnet 32.00% (96) 96.67% \$2.27
Claude3.5 Sonnet v1022 38.33% (115) $\uparrow$ 98.00% \$2.18 $\downarrow$
Table 2. Complementarity analysis of our method and baselines.
SWE-agent has the highest Apply rate due to the introduction of its execution feedback capability. For each issue, SWE-agent first constructs reproduction code to replicate the error, then verifies whether each generated patch has resolved the problem. If not, it uses feedback from the executor to iteratively refine the patch. Because it operates in a real production environment, automatically setting up and obtaining the user’s runtime environment may face challenges such as security concerns, environmental complexity, and resource constraints. Therefore, our approach focuses more on understanding the entire repository information. However, to verify the complementarity of the methods, we integrated execution feedback into LingmaAgent. The detailed results are shown in Table 2 and Figure 4. In Figure 4, we compared the issue-solving distribution of three Agent-based methods. We found that our method is highly complementary to the SWE-agent method. The two methods jointly solved 80 examples, achieving a task resolved rate of 26.67%, which further illustrates the complementarity of our method and the execution feedback method. To further integrate the execution feedback, we followed the practice and prompts of SWE-agent (Yang et al., 2024). First, we prompted the agent to write code to reproduce the problem, then fix the program and run the reproduction code to determine whether the issue is resolved. If it is not resolved, the agent debugs according to the running results and iteratively refines the generated code to improve the model’s output. The experimental result is shown as LingmaAgent (w.feedback) in Table 2, which we found to be consistent with our expectations and achieved further optimal performance (27.7%), further verifying the complementarity of LingmaAgent and SWE-agent.
Furthermore, we experimented with different models and found that performance improved significantly with more advanced models. As model capabilities continue to improve and costs decrease, we anticipate even better results in the future. Notably, Claude3.5 Sonnet v1022 (Anthropic, 2024) achieved the highest resolution rate of 38.33% while maintaining a lower average cost, demonstrating the potential for more efficient and effective software engineering problem-solving as LLMs evolve. In addition, We observed that some approaches on SWE-bench leverage voting and test-time scaling (Antoniades et al., 2024; Zhang et al., [n. d.]) to enhance performance. However, these methods may introduce significant latency in real-world applications. The exploration of efficient strategies to balance performance gains and latency in practical settings is left for future work.
@@
Figure 4. Venn diagrams of resolved cases of Lingma Agent, SWE-agent and AutoCodeRover.
4.3. Randomness in LingmaAgent
Product performance stability is crucial for user experience. However, given the inherent randomness in LLMs, it is essential to rigorously assess the consistency of our method’s outputs. Therefore, following the practices of AutoCodeRover (Zhang et al., 2024b) and SWE-agent (Yang et al., 2024), we run the system three times to evaluate its average performance and Pass@k performance. These results are shown in Table 3, where Run1-3 represents three different runs, Average represents the average performance of three times, and All represents the result of Pass@3.
Notably, LingmaAgent consistently outperforms both SWE-agent and AutoCodeRover across all three runs, with an average improvement of 18.11% over SWE-agent, the stronger baseline. The Pass@3 performance (represented by ”All” in the table 3) shows that LingmaAgent achieves a success rate of 30.67%, which is a 12.14% improvement over SWE-agent and a 37.35% improvement over AutoCodeRover, indicating its superior ability to solve problems when given multiple attempts. This suggests that the model has the potential to address a wider range of issues, and improving its pass@1 performance could be a promising approach to further enhance its issue-solving capabilities, such as sampling multiple trajectories for DPO/PPO (Rafailov et al., 2024; Wang et al., 2024b; Zheng et al., 2023a) training, which we leave for future work.
Run AutoCodeRover SWE-agent LingmaAgent
Run1 16.00% 17.33% 21.33% (23.08% $\uparrow$)
Run2 15.67% 18.00% 20.00% (11.11% $\uparrow$)
Run3 16.67% 18.00% 21.67% (20.39% $\uparrow$)
Average 16.11% 17.78% 21.00% (18.11% $\uparrow$)
All 22.33% 27.35% 30.67% (12.14% $\uparrow$)
Table 3. Performance for 3 separate runs of LingmaAgent on SWE bench Lite.
@@ 4.4.1. Module Analysis
This ablation experiment aims to study the effectiveness of LingmaAgent’s global repository understanding component. (1) Remove only the call graph module: Only the structure tree in MCTS is retained for tree search, and the reference extension module (i.e., the call relationship graph) is removed. This experiment aims to verify the effectiveness of the reference extension module, i.e., the importance of reference relations in the repository. (2) Remove only the summary module: Only the signature and dependency structure of relevant information in the repository obtained by MCTS are used as global experience, and the summary and planning of information are removed. This experiment aims to verify the effectiveness of the summary agent, i.e., the importance of comprehensive summary of repository information. (3) Remove MCTS & summary modules: LingmaAgent has no prior knowledge of the repository structure and functions, that is, it lacks empirical information about the whole repository and can only locate relevant code snippets by searching through limited information in the issue. (4) Add a review agent module: After LingmaAgent generates a patch that can be applied, in order to simulate the code review process in the development process, a static review of the patch by the review agent is added to discover possible defects in the newly generated code. If there is a defect, the patch is regenerated according to the review reason until a patch that passes the review is generated. This process is repeated up to three times.
Our experimental results demonstrate the importance of global experience and the effectiveness of the summary agent. As shown in Table 4, removing these modules all resulted in a drop in the performance of LingmaAgent, especially after removing the MCTS & summary agent; the number of problem instances solved rate decreased from 21.33% to 16.00%, which highlights the importance of global experience for automatically solving repository-level issues. In addition, we found that after adding the review agent, the performance of LingmaAgent dropped, suggesting the limitations of static review. We speculate that the LLM-based static review may only rely on the surface grammatical information of the code and cannot fully understand the semantic meaning of the code. Therefore, the static review may ignore some hidden logical errors or illogical situations in the code. Therefore, we suggest that subsequent work can combine dynamic program analysis (Zhang et al., 2023a; Deng et al., 2023; Xia et al., 2024) such as program instrumentation (Hollingsworth et al., 1994; Huang, 1978) to improve the reliability of the LLM Agent.
@@
LingmaAgent 21.33% 85.67%
- w.o. call_graph 19.67% 83.00%
- w/o. summary 17.67% 85.33%
- w/o. mcts & summary 16.00% 83.33%
- w. review 18.33% 87.67%
Table 4. Ablation results of LingmaAgent.
4.4.2. Fault Localization Analysis
Method Function_Loc File_Loc
AutoCodeRover 42.3% 62.3%
SWE-agent 45.3% 61.0%
LingmaAgent 49.3% 67.7%
Table 5. Fault localization results of LingmaAgent.
In addition to the issue resolution evaluation, we conducted a analysis focusing on the fault localization capability (Correct Location) (S et al., 2024) of LingmaAgent. The fault localization module is an intermediate module in the issue solving process. Whether the fault location is correctly located affects the subsequent program repair. Specifically, we extracted the fault locations from the developer patch and the patch generated by the model, respectively, and calculated the localization success rate by calculating whether the fault locations were consistent. This analysis aims to further illustrate the effectiveness of our repository understanding module. The resluts are shown in Table 5. We compared the two SOTA agent-based methods, AutoCodeRover and SWE-agent, where Function represents the accuracy of fault function location and File represents the accuracy of defect file location. Our findings show that our method significantly outperforms the other two methods in the success rate of fault localization at both the Function and File levels, which shows that understanding the repository and exploring critical information notably contribute to improving fault localization.